OpenStudy (anonymous):

Why does my calculator show e^-11 is .000017 when the right answer is 1.67?

OpenStudy (anonymous):

because the right answer is not 1.67

OpenStudy (mertsj):

e^-11 means "the reciprocal of e^11"

OpenStudy (anonymous):

"e^(-11) = 1.67017008 × 10-5" according to google, what is the correct answer ?

OpenStudy (anonymous):

that equates to 1/e^11 which sounds like the output of your caculator how are you getting your answer?

OpenStudy (anonymous):

10^-5 = 0.00001

OpenStudy (anonymous):

e^-11 = 0.0000167...

OpenStudy (mathmate):

There are different settings for the display. You may have selected fixed decimal display. Scientific display will give more significant figures. On mine, it shows 0.000016702, the right answer should be 1.670170079024566*10^-5 if you display in scientific notation.

OpenStudy (anonymous):

i am submitting this .000017 into my online hw and it marks it as wrong

OpenStudy (anonymous):

that is not an exact answer but a rounding

myininaya (myininaya):

does it want it a certain way?

OpenStudy (anonymous):

mathmate's answer looks right

OpenStudy (mathmate):

Online answers should specify the number of significant figures they expect, 6 decimal places, or 4 significant figures, etc.

OpenStudy (anonymous):

doesnt specify

OpenStudy (anonymous):

how do i change the display/settings to get mathmate's answer?

OpenStudy (mathmate):

Perhaps at the beginning of the instructions.

myininaya (myininaya):

can you type this 1/e^(11)

myininaya (myininaya):

\[\frac{1}{e^{11}}\]

OpenStudy (mathmate):

Good point!

OpenStudy (mathmate):

Or e^(-11)

OpenStudy (anonymous):

it becomes e^-11 then diamond enter gives .000017 and not the 1.67

myininaya (myininaya):

can you enter 1/e^(11) in your homework?

OpenStudy (anonymous):

got it right. thanks myininaya

OpenStudy (anonymous):

but still dont know why theres two diff answers for it on calc, 1.67 and .000017?

myininaya (myininaya):

it probably says to simplify and that just means for this problem it wanted you to write the number without negative exponent

myininaya (myininaya):

you said google was giving you the answer \[1.67 \cdot 10^{-5}=.0000167\]

OpenStudy (anonymous):

oh 167 rounds to 17. ohhhhhh!

myininaya (myininaya):

some calculators will give you more digits then other calculators it just depends yes!

myininaya (myininaya):

well .0000167 rounds to .000017 but i know what you mean

myininaya (myininaya):

you can round 167 to 170

OpenStudy (anonymous):

thank you

myininaya (myininaya):

np have a great night

OpenStudy (anonymous):

To 40 digits according to Mathematica: 0.00001670170079024565931263551736058087907794